OpenStudy (anonymous):

(x^3yz^-2) ---------- x^4y-x^3z How do you simplify this?

OpenStudy (anonymous):

please write question clearly show which one is the exponent etc

OpenStudy (anonymous):

\[(x ^{3}yz ^{-2})\div x ^{4}y-x ^{3}z\]

OpenStudy (id21):

\[\frac{x^3yz^{-2}}{x^4y-x^3z}=\frac{x^3yz^{-2}}{x^3(xy-z)}\] \[\frac{yz^{-2}}{xy-z}\]

OpenStudy (anonymous):

oh okay!! so do you bring the \[z ^{-2}\] to the denominator?

OpenStudy (id21):

\[\frac{y}{xyz^2-z^3}=\frac{1}{xz^{2}-\frac{z^3}{y}}\]

OpenStudy (id21):

don't know which one is simpler.

OpenStudy (anonymous):

\[x^3y^-2 \div x^4y-x^3z \]\[x^3y \div x^4 y-x^3z^3\]\[x^3y \div x^3(xy-z^3)\]\[y \div xy-z^3\]

OpenStudy (anonymous):

i dont know how to insert the line so i used divide sign hope i could help you!

OpenStudy (id21):

you can insert line using this syntax: "frac{a}{b}" try it! :)

OpenStudy (anonymous):

oh okay i get it! thanks so much! i think roh is right but im not sure..

OpenStudy (anonymous):

i am sure about my answer it is correct but again i couldnt insert the line :(

OpenStudy (id21):

Roh, You miss z^2